Three cars, car X, car Y, and car Z, begin accelerating from rest at the same time. Car X is more massive than car Y, which is more massive than car Z. the net accelerating force exerted on each car is identical. After 10 seconds, which car has the most amount of kinetic energy?

(a) They all have the same amount of kinetic energy

(b) Car X

(c) Car Y

(d) Car Z

Please explain why in your answer.
Answer #1

here,

car X is more massive than Y which more massive than Z

mX > mY > mZ

when the applied force is same

using Newton's second law

(F = m * a => a = F/m)

aX < aY < aZ

initial speed , u = 0 m/s

after same time

(from second equation of motion

s = u * t + 0.5 * a * t^2

s = 0.5 * a * t^2

for same time)

sX < sY < sZ

as the force applied is same( Work done = F * s)

the work done

Wx < Wy < Wz

using Work - energy theorm

work done = change in kinetic energy

so, the car with most kinetic energy is (d) car Z